# AI-powered Digital Twin of the Ocean: Reliable Uncertainty Quantification for Real-time Wave Height Prediction with Deep Ensemble

### v1

@@ Abstract
Environmental pollution and the depletion of fossil fuels have prompted the need for eco-friendly power generation methods based on renewable energy. However, renewable energy sources often face challenges in providing stable power due to low energy density and non-stationary. Wave energy converters (WECs), in particular, need reliable real-time wave height prediction to address these issues caused by irregular wave patterns, which can lead to the inefficient and unstable operation of WECs. In this study, we propose an AI-powered reliable real-time wave height prediction model, aiming both high predictive accuracy and reliable uncertainty quantification (UQ). The proposed architecture LSTM-DE, integrates long short-term memory (LSTM) networks for temporal prediction with deep ensemble (DE) for robust UQ, achieving accuracy and reliability in wave height prediction. To further enhance the reliability of the predictive models, uncertainty calibration is applied, which has proven to significantly improve the quality of the quantified uncertainty. Based on the real operational data obtained from an oscillating water column-wave energy converter (OWC-WEC) system in Jeju, South Korea, we demonstrate that the proposed LSTM-DE model architecture achieves notable predictive accuracy ($R^{2}>0.9$) while increasing the uncertainty quality by over 50% through simple calibration technique. Furthermore, a comprehensive parametric study is conducted to explore the effects of key model hyperparameters, offering valuable guidelines for diverse operational scenarios, characterized by differences in wavelength, amplitude, and period. The findings show that the proposed method provides robust and reliable real-time wave height predictions, facilitating digital twin of the ocean.
@@ 1 Introduction
Energy decarbonization and sustainable industries are essential for addressing global climate change, as emphasized in the sustainable development goals (SDGs) adopted by all United Nations Member States in 2015 [1, 2]. These goals highlight the transformation from conventional fossil fuels to renewable energy resources, which are capable of reducing greenhouse gas emissions and mitigating climate change. From this perspective, the utilization of renewable energy resources and technologies is continuously required. The most widely used renewable energy types are wind energy, solar energy, and hydropower. Especially, ocean waves generated by wind, with their large potential energy capacity and higher energy density compared to other renewable resources [3], are expected to play a significant role in achieving SDG 7 (Affordable and Clean Energy) and SDG 13 (Climate Action). Consequently, global interest in wave energy-based power generation has been steadily increasing over time [4].
@@
In perspective of wave energy conversion technology, various systems called WECs have been proposed and have experienced rapid growth in recent decades to utilize the enormous energy resources in the ocean [5, 6, 7, 8]. Primary wave energy conversion is achieved by oscillating water column-based systems, including floating body, oscillating solid member, and oscillating water within a structure [9, 10, 11]. Unlike hydropower energy, the power generation using WECs mainly depends on short-term and local conditions such as wave height, period, and spectra. It means that these systems may face several limitations in stable power generation due to local and irregular wave patterns. Therefore, real-time wave height prediction is essential to increase the availability on WECs. In fact, recent studies have demonstrated the importance of real-time wave height prediction model for the operation of the WECs [12, 13, 14, 15, 16]. Moreover, these prediction models are also necessary to construct digital twins of the WECs, which are virtual representations of physical objects and systems, to make real-time decisions regarding the stable operation.
@@
There have been numerous studies on wave height prediction, focusing on the development of theoretical and statistical models [17, 18, 19, 20, 21, 22, 23]. More recently, machine learning (ML)-based approaches, including deep learning (DL), have been increasingly used to overcome the limitations of conventional methodologies, particularly in handling nonlinear patterns and large datasets [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. These studies involve various datasets provided by simulation, experiment, operational, and public data repositories, covering scales from local to global. An oscillating water column-wave energy converter (OWC-WEC), which is the primary focus of this study, also requires high-quality and reliable wave height prediction models for its operation and management. Various studies have explored numerical and experimental approaches [37, 38, 39, 40], as well as artificial intelligence (AI)-based approach using the LSTM algorithm to handle time-series wave patterns data [41]. However, most existing wave height prediction models are limited in that they cannot provide the information on the uncertainty over their predictions. More specifically, their deterministic models provide a single prediction value, while probabilistic models provide not only the prediction values, but also the quantified uncertainty in their predictions. In real-world applications, deterministic prediction models cannot provide uncertainty estimates, which may lead to catastrophic outcomes in WEC operations, such as excessive air inflow resulting from unanticipated changes in wave height that the deterministic model fails to account for, potentially damaging the air turbine.
@@
From a DL perspective on UQ, there are two main categories [42]: Bayesian-based approach [43, 44] and ensemble-based approach [45]. The Bayesian-based approach, commonly known as Bayesian neural network (BNN) has been widely used for uncertainty quantification (UQ) in DL models [46]. For real-time ocean wave prediction, a BNN-based prediction method was proposed in [47]. On the other hand, the ensemble-based approach, known as DE, is also often used to effectively calculate uncertainty owing to their advantages: simplicity and scalability. In particular, DE is considered more practical than BNN for dataset shifting, as it explores diverse modes in function space, whereas Bayesian-based approach tends to fall into a single mode [48]. Despite these advantages, there are no known applications of DE in wave height prediction, especially for short-term and local scales, although research has been conducted on probabilistic spatiotemporal forecasting for solar irradiation on a global scale [49]. Furthermore, in uncertainty estimation for AI-based classification and regression models, including DL, it is crucial to conduct quantitative quality evaluations to ensure the reliability of the probabilistic model and apply uncertainty calibrations based on these evaluations metrics [50, 51, 52, 53, 54, 55]. However, prior studies, including the BNN-based method for wave height prediction [47], have not incorporated calibration methods to enhance the quality of UQ. In this regard, the calibration framework for DEs have been proposed in multi-output regression tasks [56] to obtain reliable probabilistic predictions. Although this research represents the state-of-the-art reliable probabilistic predictions in multi-output regression tasks, which could be applicable to time-series data like wave patterns, it has not yet been validated for temporal multi-output regression tasks.
@@
In this study, we propose a simple and scalable LSTM-DE model architecture to construct a reliable real-time wave height prediction model to be embedded into digital twin of WECs. Specifically, we conduct the research mainly considering the following three aspects: (1) a simplicity and scalability of LSTM-DE model architecture, (2) high accuracy for prediction, and (3) reliable UQ coupled with uncertainty calibration. Within these perspectives, we also demonstrate the impact of ocean-related domain knowledge on model performance to implement effective and efficient experimental design and optimize the hyperparameters of DL model architectures. To verify the effectiveness and applicability of embedding into digital twin of WECs, we use real operational data from the 500kw OWC-WEC installed on the coast of Jeju Island, South Korea (Fig. 1), along with its digital twin platform developed by Korea Research Institute of Ships & Ocean Engineering (KRISO) (Fig. 2). To the best of our knowledge, this is the first case of using a calibrated DE methodology for real-time wave height prediction. The corresponding results are expected to enhance the availability of WECs and facilitate the development of a digital twin of the ocean, providing robust and reliable real-time wave height predictions in perspective of temporal multi-output regression tasks.
@@ 4.1 Details of experiments
In this study, the real operational pressure data measured by the 500kw OWC-WEC shown which is installed on the coast of Jeju Island, South Korea in Fig. 1, was utilized to analyze the irregular ocean wave height variations within the chamber. The data was collected over a specific period from December 2020 to February 2021, capturing pressure fluctuations at the bottom of the OWC-WEC chamber with a sampling frequency of 20Hz (0.05 seconds for 1 timestep). It focused primarily on winter conditions, which are expected to provide suitable operational scenarios for wave energy conversion. As mentioned in Section 2.1, these pressure values, initially recorded in millibars (mbar), were converted to meters (m) to predict wave heights on a meter scale, based on the principles of hydrostatic pressures. Additionally, min-max normalization was applied to appropriately scale the data. This dataset was split into train (80%), validation (10%), and test (10%) based on time sequence, as illustrated in Fig. 8. For UQ calibration, a post-processing step with STD scaling, a separate validation dataset independent of the training and test process, is required [55]. Therefore, the validation dataset is specifically utilized for this purpose.

### v2

@@ Abstract
Environmental pollution and fossil fuel depletion have prompted the need for renewable energy-based power generation. However, its stability is often challenged by low energy density and non-stationary conditions. Wave energy converters (WECs), in particular, need reliable real-time wave height prediction to address these issues caused by irregular wave patterns, which can lead to the inefficient and unstable operation of WECs. In this study, we propose an AI-powered reliable real-time wave height prediction model that integrates long short-term memory (LSTM) networks for temporal prediction with deep ensemble (DE) for robust uncertainty quantification (UQ), ensuring high accuracy and reliability. To further enhance the reliability, uncertainty calibration is applied, which has proven to significantly improve the quality of the quantified uncertainty. Using real operational data from an oscillating water column-wave energy converter (OWC-WEC) system in Jeju, South Korea, the model achieves notable accuracy ($R^{2}>0.9$), while increasing uncertainty quality by over 50% through simple calibration technique. Furthermore, a comprehensive parametric study is conducted to explore the effects of key model hyperparameters, offering valuable guidelines for diverse operational scenarios, characterized by differences in wavelength, amplitude, and period. These results demonstrate the model’s capability to deliver reliable predictions, facilitating digital twin of the ocean.
@@ 1 Introduction
Energy decarbonization and sustainable industries are essential for addressing global climate change, as emphasized in the sustainable development goals (SDGs) adopted by all United Nations Member States in 2015 [1, 2]. These goals highlight the transformation from conventional fossil fuels to renewable energy resources, which are capable of reducing greenhouse gas emissions and mitigating climate change. From this perspective, the utilization of renewable energy resources and technologies is continuously required. The most widely used renewable energy types are wind energy, solar energy, and hydropower. Especially, ocean waves generated by wind, with their large potential energy capacity and higher energy density compared to other renewable resources [3], are expected to play a significant role in achieving SDG 7 (Affordable and Clean Energy) and SDG 13 (Climate Action). Consequently, global interest in wave energy-based power generation has been steadily increasing over time [4, 5].
@@
In perspective of wave energy conversion technology, various systems called WECs have been proposed and have experienced rapid growth in recent decades to utilize the enormous energy resources in the ocean [6, 7, 8, 9]. Primary wave energy conversion is achieved by oscillating water column-based systems, including floating body, oscillating solid member, and oscillating water within a structure [10, 11, 12]. Unlike hydropower energy, the power generation using WECs mainly depends on short-term and local conditions such as wave height, period, and spectra. It means that these systems may face several limitations in stable power generation due to local and irregular wave patterns. Therefore, real-time wave height prediction is essential to increase the availability on WECs. In fact, recent studies have demonstrated the importance of real-time wave height prediction model for the operation of the WECs [13, 14, 15, 16, 17, 18]. Moreover, these prediction models are also necessary to construct digital twins of the WECs, which are virtual representations of physical objects and systems, to make real-time decisions regarding the stable operation.
@@
There have been numerous studies on wave height prediction, focusing on the development of theoretical and statistical models [19, 20, 21, 22, 23, 24, 25]. More recently, machine learning (ML)-based approaches, including deep learning (DL), have been increasingly used to overcome the limitations of conventional methodologies, particularly in handling nonlinear patterns and large datasets [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. These studies involve various datasets provided by simulation, experiment, operational, and public data repositories, covering scales from local to global. An oscillating water column-wave energy converter (OWC-WEC), which is the primary focus of this study, also requires high-quality and reliable wave height prediction models for its operation and management. Various studies have explored numerical and experimental approaches [41, 42, 43, 44, 45], as well as artificial intelligence (AI)-based approach using the LSTM algorithm to handle time-series wave patterns data [46]. However, most existing wave height prediction models are limited in that they cannot provide the information on the uncertainty over their predictions. More specifically, their deterministic models provide a single prediction value, while probabilistic models provide not only the prediction values, but also the quantified uncertainty in their predictions. In real-world applications, deterministic prediction models cannot provide uncertainty estimates, which may lead to catastrophic outcomes in WEC operations, such as excessive air inflow resulting from unanticipated changes in wave height that the deterministic model fails to account for, potentially damaging the air turbine.
@@
From a DL perspective on UQ, there are two main categories [47]: Bayesian-based approach [48, 49] and ensemble-based approach [50]. The Bayesian-based approach, commonly known as Bayesian neural network (BNN) has been widely used for UQ in DL models [51]. For real-time ocean wave prediction, a BNN-based prediction method was proposed in [52]. On the other hand, the ensemble-based approach, known as DE, is also often used to effectively calculate uncertainty owing to their advantages: simplicity and scalability. In particular, DE is considered more practical than BNN for dataset shifting, as it explores diverse modes in function space, whereas Bayesian-based approach tends to fall into a single mode [53]. Despite these advantages, there are no known applications of DE in wave height prediction, especially for short-term and local scales, although research has been conducted on probabilistic spatiotemporal forecasting for solar irradiation on a global scale [54]. Furthermore, in uncertainty estimation for AI-based classification and regression models, including DL, it is crucial to conduct quantitative quality evaluations to ensure the reliability of the probabilistic model and apply uncertainty calibrations based on these evaluations metrics [55, 56, 57, 58, 59, 60]. However, prior studies, including the BNN-based method for wave height prediction [52], have not incorporated calibration methods to enhance the quality of UQ. In this regard, the calibration framework for DEs have been proposed in multi-output regression tasks [61] to obtain reliable probabilistic predictions. Although this research represents the state-of-the-art reliable probabilistic predictions in multi-output regression tasks, which could be applicable to time-series data like wave patterns, it has not yet been validated for temporal multi-output regression tasks.
@@
In this study, we propose a simple and scalable LSTM-DE model architecture to construct a reliable real-time wave height prediction model to be embedded into digital twin of WECs. Specifically, we conduct the research mainly considering the following three aspects: (1) a simplicity and scalability of LSTM-DE model architecture, (2) high accuracy for prediction, and (3) reliable UQ coupled with uncertainty calibration. Within these perspectives, we also demonstrate the impact of ocean-related domain knowledge on model performance to implement effective and efficient experimental design and optimize the hyperparameters of DL model architectures. To verify the effectiveness and applicability of embedding into digital twin of WECs, we use real operational data from the 500kw OWC-WEC installed on the coast of Jeju Island, South Korea (Fig. 1) [62], along with its digital twin platform developed by Korea Research Institute of Ships & Ocean Engineering (KRISO) (Fig. 2). To the best of our knowledge, this is the first case of using a calibrated DE methodology for real-time wave height prediction. The corresponding results are expected to enhance the availability of WECs and facilitate the development of a digital twin of the ocean, providing robust and reliable real-time wave height predictions in perspective of temporal multi-output regression tasks.
@@ 4.1 Details of experiments
In this study, the real operational pressure data measured by the 500kw OWC-WEC shown which is installed on the coast of Jeju Island, South Korea in Fig. 1, was utilized to analyze the irregular ocean wave height variations within the chamber. The data was collected over a specific period from December 2020 to February 2021, capturing pressure fluctuations at the bottom of the OWC-WEC chamber with a sampling frequency of 20Hz (0.05 seconds for 1 timestep). It focused primarily on winter conditions, which are expected to provide suitable operational scenarios for wave energy conversion [74]. As mentioned in Section 2.1, these pressure values, initially recorded in millibars (mbar), were converted to meters (m) to predict wave heights on a meter scale, based on the principles of hydrostatic pressures. Additionally, min-max normalization was applied to appropriately scale the data. This dataset was split into train (80%), validation (10%), and test (10%) based on time sequence, as illustrated in Fig. 8. For UQ calibration, a post-processing step with STD scaling, a separate validation dataset independent of the training and test process, is required [60]. Therefore, the validation dataset is specifically utilized for this purpose.